OpenStudy (anonymous):

Solve for : C = 2 r.

OpenStudy (anonymous):

solve what?

OpenStudy (anonymous):

\[\pi\]

OpenStudy (anonymous):

I believe C is the circumference C = 2 pi r pi = C/(2r)

OpenStudy (ash2326):

Circumference =2pi*r c=2pi*r pi=C/2r